# QT-TDM: Planning with Transformer Dynamics Model and Autoregressive Q-Learning

### v1

@@ I INTRODUCTION
Transformer dynamics models (TDMs) have proven effective in background planning scenarios, where an actor-critic model is trained on the imagined trajectories generated by the learned dynamics model. During inference, the learned actor-critic model selects the suitable actions. TDMs show an outstanding performance in discrete action spaces [12], [13] and in long-term memory tasks [14].
In real-time planning scenarios, where the learned dynamics model plans ahead by being unrolled forward from the current state to select the best action, TDMs encounter hurdles. Specifically, inference is slow and computationally inefficient [15] due to the autoregressive token prediction and the per-dimension tokenization scheme, which increases sequence length as the environment’s dimensionality increases. This makes planning for long horizons impractical, especially in the robotics domain, where fast inference is essential. Therefore, TDMs require more optimization on the architecture level, and more sample-efficient planning algorithms are needed to achieve faster real-time inference.
Figure 1: QT-TDM Inference. The learned TDM model plans for short planning horizon $H$, while the learned QT model estimates an autoregressive terminal value $Q^{i}_{H}$ for each action dimension $a^{i}_{H}$ which guides the planning beyond the myopic horizon.
To this end, we introduce QT-TDM, a model-based algorithm that combines the strengths of a TDM and a model-free Q-Transformer (QT) [16]. Inspired by the TD-MPC algorithm [17], our proposed model achieves fast inference (as shown in Fig. 1) by combining a short planning horizon with a terminal value that is estimated by the Q-Transformer model which provides an estimation of a long-term return beyond the myopic planning horizon. Additionally, the sequence length is reduced by tokenizing the high-dimensional state space into a single token using a learned linear layer [11], as opposed to the conventional per-dimension tokenization method [15], [18].
@@ II Related Work
Transformers in Reinforcement Learning. Following recent advancements across various domains driven by Transformers, reinforcement learning has emerged as the next significant challenge for Transformers to tackle. The key idea is to reformulate the RL problem into sequence modeling problem. The Decision Transformer model (DT) [11] represents the RL problem as conditional sequence modeling. By conditioning on desired target return values, DT can autoregressively generate the desired trajectories. Concurrently, the Trajectory Transformer model (TT) [18] utilizes beam search as a planning algorithm during execution. TT performs well on long-horizon tasks and can perform imitation learning and goal-conditioned RL under the same framework. The Gated Transformer-XL (GTrXL) [21] stabilizes the learning of standard Transformer with RL objectives by replacing residual connections with gating layers and introducing Identity Map Reordering. GTrXL surpasses LSTMs on memory-based environments. The Q-Transformer model (QT) [16] proposes an autoregressive Q-learning (see section III) that predicts a Q-value for each action dimension, after discretizing each action dimension individually into uniformally-spaced bins. QT outperforms DT and other state-of-the-art models in large-scale robotic offline RL tasks.
@@
Transformer Dynamics Model. Motivated by the success of Transformers in sequence modeling tasks, there has been a lot of recent attention on using Transformers as dynamics models. One of the earliest attempts is TransDreamer [14] which as implied by the name is a modification of the Dreamer model [22]. TransDreamer replaces the Recurrent State-Space Model (RSSM) [23] with a Transformer State-Space Model (TSSM), improving TransDreamer’s performance in long-term memory tasks. IRIS [12] and TWM [13] are two sample-efficient model-based agents that are trained inside the imagination of a Transformer-based world model. IRIS’ world model consists of a discrete autoencoder [24] as an observation model and a GPT-like Transformer [8] as a dynamics model, while the world model of TWM consists of a variational autoencoder [25] and a Transformer-XL [9]. Both models work with discrete action environments and they achieve impressive results on the Atari 100K benchmark.
Generalist TDM [15] is the first attempt to use a learned TDM for continuous real-time planning with Model Predictive Control. Generalist TDM performs well in a single environment (i.e, specialist setting) and generalizes to unseen environments (i.e., generalist setting), in a few-shot and in zero-shot scenarios. Despite of its capabilities, it has two shortcomings. First, the training data is collected by an expert agent and not by its own interactions with the environment. Second, it suffers from slow inference because of the long-horizon planning and because of the design choices that are based on the Gato Transformer model [26] which uses the per-dimension tokenization scheme.
@@
To overcome the above shortcomings, we introduce the QT-TDM model, which explores the environment to collect training data and has faster inference speed by shortening the planning horizon and utilizing the Q-Transformer model [16] to estimate a long-term return beyond the short-term planning horizon.
@@ III Background
Figure 2: QT-TDM Architecture, which consists of two modules: (a) TDM and (b) QT. Both modules have a GPT-like Transformer as a main component and share the same tokenization scheme. The state $s_{t}$ is tokenized into a single token using a learned linear layer. A per-dimension tokenization is performed for the $N$-dimensional action by discretizing each dimension independently into $K$ bins, then using an embedding table. The TDM module predicts the next state $\hat{s}_{t+1}$ and the reward $\hat{r}_{t}$ and is trained on $\mathit{L}$ sampled time steps (for brevity, we only show two time steps). The QT module predicts a Q-value for each action dimension $\hat{q}_{t}^{i,1:K}$.
@@
Reinforcement Learning. We formulate the problem of continuous control as an infinite-horizon Markov Decision Process (MDP) that can be formalized by a tuple $(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{T},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the continuous action space, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is a reward function, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ is the transition function, and $\gamma\in[0,1]$ is a discount factor. The goal of reinforcement learning is to learn a policy $\pi:\mathcal{S}\mapsto\mathcal{A}$ from interacting with the environment that maximizes the expected cumulative discounted reward $\mathbb{E}_{\pi}[\sum^{\infty}_{t=0}\gamma^{t}r_{t}]$. In this work, the policy $\pi$ is derived from planning with a learned model of the environment’s dynamics.
@@
Autoregressive Q-Learning. Applying Q-learning with Transformers is challenging since Transformers require discretizing the action space into tokens to effectively apply the attention mechanism. Therefore, the standard Q-learning needs to be reformulated in order to be applied. In the Q-Transformer model [16], an autoregressive Q-learning formulation is proposed where each action dimension is treated as a separate time step. This way, each action dimension can be discretized individually, rather than discretizing the entire action space, thus avoiding the exponential growth in the discrete action space. An autoregressive discrete Q-function is employed which predicts a separate Q-value for each action dimension.
Let $a_{t}=(a_{t}^{1},...,a_{t}^{N})$ be an $N$-dimensional action at time step $t$. The autoregressive Q-function predicts a Q-value for each action dimension $a_{t}^{i}$ that is conditioned on the state $s_{t}$ and the previous action dimensions $a_{t}^{1:i-1}$:
@@
The Q-targets of all action dimensions except the last one are computed by maximizing over the discretized bins of their subsequent dimension within the same time step. The Q-target of the last dimension is computed by the discounted maximization of the first dimension of the next time step plus the reward. The reward is only applied on the last dimension as it is observed after executing the whole action. In addition, the Q-values are only discounted between time steps (i.e., discount factor $\gamma$ is only applied for the last dimension), to ensure the same discounting as in the original MDP.
The Q-Transformer model [16] has been evaluated in an offline RL fashion with large-scale robotic sparse reward tasks. In this work, we utilize the Q-Transformer model in an online RL fashion to estimate a terminal Q-value in a short-horizon real-time planning task, in order to achieve faster planning.
IV Method
@@ IV-A Architecture
The dynamics model is implemented as a GPT-like Transformer [8] that computes a deterministic hidden state $h_{t}$ based on the history of past steps. We only consider the hidden state that corresponds to the last action dimension (see Fig. 2(a); orange boxes vs. faded gray boxes). Predictors for the next state and reward are conditioned on the hidden state which are implemented as multilayer perceptrons (MLPs). The model components are as follows:
@@
where the reward model outputs the mean of a normal distribution.
The Q-Transformer model consists of a GPT-like Transformer and an autoregressive discrete Q-function that predicts a Q-value for each action dimension which is implemented as MLP. The Transformer computes a deterministic hidden state $h_{t}^{i}$ for each action dimension $a_{t}^{i}$ based on the state $s_{t}$ and previous action dimensions $a_{t}^{1:i-1}$. The model components are as follows:
@@
Both models, TDM and QT, tokenize the input sequences in the same way. Let $s\in\mathcal{S}$ be an $M$-dimensional state and $a\in\mathcal{A}$ is an $N$-dimensional continuous action. We follow [11] in tokenizing the state $s$ into a single token obtained with a learned linear layer, rather than the conventional per-dimension tokenization [15], [18] which increases the input sequence length. We perform a per-dimension tokenization for the $N$-dimensional continuous action $a=(a^{1},a^{2},....,a^{N})$ by discretizing each dimension independently into $K$ uniformally-spaced bins, then invoking the token embedding from a learned embedding table. TDM takes as input a sequence of $\mathit{L}(N+1)$ tokens, where $\mathit{L}$ is time steps. QT takes as input a sequence of $N$ tokens as it ignores the last action dimension.
Algorithm 1 QT-TDM (training)
1:$\theta$: initialized TDM parameters
@@
2:for each training step do // Collect episode with QT-TDM and add to buffer
3:     $\mathcal{B}\leftarrow\mathcal{B}\cup\{s_{t},a_{t},r_{t},s_{t+1}\}_{t=0}^{T-1}$
4:     for num updates per episode do
5:         $\{s_{t},a_{t},r_{t},s_{t+1}\}_{t=1}^{L}\sim\mathcal{B}$$\triangleright$ Sample traj.
@@
6:         for $t=1...L$ do
7:              $h_{t}=\mathit{f}_{\theta}(s\leq t,a^{1:N}\leq t)$$\triangleright$ Hidden state.
8:              $\hat{s}_{t+1}=\mathit{g}_{\theta}(h_{t})$$\triangleright$ Transition.
9:              $\hat{r}_{t}\sim\mathit{p}_{\theta}(\hat{r}_{t}|h_{t})$$\triangleright$ Reward.
10:         $\theta\leftarrow\theta-\eta_{d}\gradient_{\theta}\mathcal{L}_{\theta}^{Dyn.}$$\triangleright$ Equation 8
@@
11:         for $i=1...N$ do
12:              $h_{t}^{i}=\mathit{f}_{\phi}(s_{t},a_{t}^{1:i-1})$$\triangleright$ Hidden state.
13:              $\hat{q}_{t}^{i,1:K}=\mathit{g}_{\phi}(h_{t}^{i})$$\triangleright$ Q-Values.
14:         $\phi\leftarrow\phi-\eta_{q}\gradient_{\phi}\mathcal{L}_{\phi}^{Q}$$\triangleright$ Equation 9
@@
15:         $\bar{\phi}\leftarrow(1-\zeta)\bar{\phi}+\zeta\phi$
@@ IV-B Training
$\mathcal{L}_{\theta}^{Dyn.}=\sum^{\mathit{L}}_{t=1}\Bigl{[}\mathcal{\beta}_{1}\norm{\mathit{g}_{\theta}(h_{t})-s_{t+1}}^{2}_{2}-\mathcal{\beta}_{2}\ln{\mathit{p}_{\theta}(r_{t}|h_{t})}\Bigr{]},$ (8)
@@
Algorithm 2 QT-TDM (planning)
1:$\theta$, $\phi$: TDM parameters, QT parameters
@@
$s_{t}$, $H$, $\mathcal{I}$: current state, len. of horizon, num. of iterations
2:for $n=1...\mathcal{I}$ do
3:     Sample $J$ action seq. from $\mathcal{N}(\mu^{n-1},(\sigma^{n-1})^{2}\mathbf{I})$
4:     Sample $J_{QT}$ action seq. using QT and TDM
     // Generate trajectories using rollouts from TDM
        and estimate the total return $\mathcal{F}_{\varGamma}$
5:     for all $J+J_{QT}$ action sequences do
6:         for $t=0...H-1$ do
7:              $h_{t}=\mathit{f}_{\theta}(s\leq t,a^{1:N}\leq t)$$\triangleright$ Hidden state.
8:              $\mathcal{F}_{\varGamma}=\mathcal{F}_{\varGamma}+\gamma^{t}\mathit{p}_{\theta}(\hat{r}_{t}|h_{t})$$\triangleright$ Reward.
9:              $\hat{s}_{t+1}=\mathit{g}_{\theta}(h_{t})$$\triangleright$ Transition.              // Estimate the terminal Q-value using QT       and add the value of last action dim. to $\mathcal{F}_{\varGamma}$
10:         $\mathcal{F}_{\varGamma}=\mathcal{F}_{\varGamma}+\gamma^{H}\max\limits_{a_{H}^{N}}Q_{\phi}(\hat{s}_{H},a_{H}^{N})$
11:     Update $\mu^{n}$ and $\sigma^{n}$ using top-$k$ trajectories.
12:return $a\sim\mathcal{N}(\mu^{\mathcal{I}},(\sigma^{\mathcal{I}})^{2}\mathbf{I})$

### v2

@@ Abstract
Index Terms:
Model learning for control, machine learning for robot control, deep learning methods.
@@ I Introduction
Transformer dynamics models (TDMs) [12], [13] have proven effective in background planning [14] scenarios, where an actor-critic model is trained on the imagined trajectories generated by the learned dynamics model. During inference, the learned actor-critic model selects the suitable actions. TDMs show an outstanding performance in discrete action spaces [13], [15] and in long-term memory tasks [16].
In real-time planning scenarios, where the learned dynamics model plans ahead by being unrolled forward from the current state to select the best action, TDMs encounter hurdles. Specifically, inference is slow and computationally inefficient [12], [17] due to the autoregressive token prediction and the per-dimension tokenization scheme, which increases sequence length as the environment’s dimensionality increases. This makes planning for long horizons impractical, especially in the robotics domain, where fast inference is essential. Therefore, TDMs require more optimization on the architecture level, and more sample-efficient planning algorithms are needed to achieve faster real-time inference.
Figure 1: QT-TDM Inference: The learned TDM model plans for short planning horizon $H$, while the learned QT model estimates an autoregressive terminal value $Q^{i}_{H}$ for each action dimension $a^{i}_{H}$ which guides the planning beyond the myopic horizon.
To this end, we introduce QT-TDM, a model-based algorithm that combines the strengths of a TDM and a model-free Q-Transformer (QT) [18]. Inspired by the TD-MPC algorithm [19], our proposed model achieves fast inference (as shown in Fig. 1) by combining a short planning horizon with a terminal value that is estimated by the Q-Transformer model which provides an estimate of a long-term return beyond the myopic planning horizon. Additionally, the sequence length is reduced by tokenizing the high-dimensional state space into a single token using a learned linear layer [11], as opposed to the conventional per-dimension tokenization method [12], [17].
@@ II Related Work
II-1 Transformer Dynamics Model
@@
Motivated by the success of Transformers in sequence modeling tasks, there has been a lot of recent attention on using Transformers as dynamics models. One of the earliest attempts is TransDreamer [16] which as implied by the name is a modification of the Dreamer model [14]. TransDreamer replaces the Recurrent State-Space Model (RSSM) [22] with a Transformer State-Space Model (TSSM), improving TransDreamer’s performance in long-term memory tasks. IRIS [13] and TWM [15] are two sample-efficient model-based agents that are trained inside the imagination of a Transformer-based world model. IRIS’ world model consists of a discrete autoencoder [23] as an observation model and a GPT-like Transformer [8] as a dynamics model, while the world model of TWM consists of a variational autoencoder [24] and a Transformer-XL [9]. Both models work with discrete action environments and they achieve impressive results on the Atari 100K benchmark. Generalist TDM [12] is the first attempt to use a learned TDM for continuous real-time planning with Model Predictive Control. Generalist TDM performs well in a single environment (i.e, specialist setting) and generalizes to unseen environments (i.e., generalist setting), in a few-shot and in zero-shot scenarios. Despite of its capabilities, it has two shortcomings. First, the training data is collected by an expert agent and not by its own interactions with the environment. Second, it suffers from slow inference because of the long-horizon planning and because of the design choices that are based on the Gato Transformer model [25] which uses the per-dimension tokenization scheme.
@@
To overcome the above shortcomings, we introduce the QT-TDM model, which explores the environment to collect training data and has faster inference speed by shortening the planning horizon and utilizing the QT model [18] to estimate a long-term return beyond the short-term planning horizon.
II-2 Robotics Foundation Models
Inspired by the success of Vision/Language Foundation Models [26], there remains significant potential for the development of specialized Robotics Foundation Models (RFMs). Foundation Models, primarily based on Transformer architectures, are pre-trained on large-scale datasets and exhibit remarkable zero-shot and few-shot generalization capabilities. Examples of RFMs are RT-2 [27], Q-Transformer [18], Gato [25] and PaLM-E [28]. All existing RFMs adopt a model-free (model-agnostic) approach. However, many researchers argue that a model-based approach based on Foundation World Models (FWMs) is a promising direction for addressing complex robotics challenges [29]. While Generalist TDM [12] shows the potential of this direction, this work builds on it and further improves real-time planning capabilities and efficiency. In the future work section, we propose strategies to further advance QT-TDM toward the realization of FWMs.
@@ III Background
Figure 2: QT-TDM Architecture, which consists of two modules: (a) TDM and (b) QT. Both modules have a GPT-like Transformer as a main component and share the same tokenization scheme. The state $s_{t}$ is tokenized into a single token using a learned linear layer. A per-dimension tokenization is performed for the $N$-dimensional action by discretizing each dimension independently into $K$ bins, then using an embedding table. The TDM module predicts the next state $\hat{s}_{t+1}$ and the reward $\hat{r}_{t}$ and is trained on $\mathit{L}$ sampled time steps (for brevity, we only show two time steps). The QT module predicts a Q-value for each action dimension $\hat{q}_{t}^{i,1:K}\hskip 5.0pt\forall i\in[1,...,N]$.
III-1 Reinforcement Learning
@@
We formulate the problem of continuous control as an infinite-horizon Markov Decision Process (MDP) that can be formalized by a tuple $(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{T},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the continuous action space, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$ is a reward function, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ is the transition function, and $\gamma\in[0,1]$ is a discount factor. The goal of reinforcement learning is to learn a policy $\Pi_{\theta}:\mathcal{S}\mapsto\mathcal{A}$ from interacting with the environment that maximizes the expected cumulative discounted reward $\mathbb{E}_{\Pi_{\theta}}[\sum^{\infty}_{t=0}\gamma^{t}r_{t}]$. In this work, the policy $\Pi_{\theta}$ is derived from planning with a learned dynamics model.
III-2 Model Predictive Control
In control, learning $\Pi_{\theta}$ is formulated as a trajectory optimization problem, in which at each step $t$, optimal actions $a_{t:t+H}$ over a finite horizon $H$ are estimated to maximize the discounted sum of rewards:
$\Pi_{\theta}(s_{t})=\arg\max_{a_{t:t+H}}\mathbb{E}\left[\sum_{i=t}^{H}\gamma^{i}r_{i}\right],$ (1)
and the first action $a_{t}$ is executed. This method is known as Model Predictive Control (MPC). Eq. 1 is not predicting long-term rewards beyond $H$. Consequently, incorporating a value function of the terminal state $s_{t+H}$ provides an estimate of the long-term return, a method referred to as MPC with a terminal value [19]. An alternative approach, known as MPC with value summation [4], involves summing value functions over a finite horizon rather than summing rewards. In this work, we utilize Q-Transformer to estimate a terminal Q-value in a myopic planning horizon.
III-3 Autoregressive Q-Learning
@@
Applying Q-learning with Transformers is challenging since Transformers require discretizing the action space into tokens to effectively apply the attention mechanism. Therefore, the standard Q-learning needs to be reformulated in order to be applied. In the Q-Transformer model [18], an autoregressive Q-learning formulation is proposed where each action dimension is treated as a separate time step. This way, each action dimension can be discretized individually, rather than discretizing the entire action space, thus avoiding the exponential growth in the discrete action space. An autoregressive discrete Q-function is employed which predicts a separate Q-value for each action dimension. Let $a_{t}=(a_{t}^{1},...,a_{t}^{N})$ be an $N$-dimensional action at time step $t$. The autoregressive Q-function predicts a Q-value for each action dimension $a_{t}^{i}$ that is conditioned on the state $s_{t}$ and the previous action dimensions $a_{t}^{1:i-1}$:
@@
The Q-targets of all action dimensions except the last one are computed by maximizing over the discretized bins of their subsequent dimension within the same time step. The Q-target of the last dimension is computed by the discounted maximization of the first dimension of the next time step plus the reward. The reward is only applied on the last dimension as it is observed after executing the whole action. In addition, the Q-values are only discounted between time steps (i.e., discount factor $\gamma$ is only applied for the last dimension), to ensure the same discounting as in the original MDP. The Q-Transformer model [18] has been evaluated in an offline RL fashion with large-scale robotic sparse reward tasks. In this work, we utilize the Q-Transformer model in an online RL fashion to estimate a terminal Q-value in a short-horizon real-time planning task, in order to achieve faster planning.
IV Methodology
@@ IV-A Architecture
TDM is implemented as a GPT-like Transformer [8] that computes a deterministic hidden state $h_{t}$ conditioned on the states and actions of past steps. We consider only the hidden state corresponding to the last action dimension, as it attends to all preceding action dimensions (see Fig. 2a; orange boxes vs. faded gray boxes). Predictors for the next state and reward are conditioned on the hidden state which are implemented as multilayer perceptrons (MLPs). The model components are as follows:
@@
the reward model outputs the mean of a normal distribution.
The Q-Transformer model consists of a GPT-like Transformer and an autoregressive discrete Q-function that predicts a Q-value for each action dimension which is implemented as MLP. The Transformer computes a deterministic hidden state $h_{t}^{i}$ for each action dimension $a_{t}^{i}$ conditioned on the state $s_{t}$ and previous action dimensions $a_{t}^{1:i-1}$. The model components are as follows:
@@
Both models, TDM and QT, tokenize the input sequences in the same way. Let $s\in\mathcal{S}$ be an $M$-dimensional state and $a\in\mathcal{A}$ is an $N$-dimensional continuous action. We follow [11] in tokenizing the state $s$ into a single token obtained with a learned linear layer, rather than the conventional per-dimension tokenization [12], [17] which increases the input sequence length. We perform a per-dimension tokenization for the $N$-dimensional continuous action $a=(a^{1},a^{2},....,a^{N})$ by discretizing each dimension independently into $K$ uniformally-spaced bins, then invoking the token embedding from a learned embedding table. TDM takes as input a sequence of $\mathit{L}\times(N+1)$ tokens, where $\mathit{L}$ is time steps. QT takes as input a sequence of $N$ tokens as it ignores the last action dimension.
Require: $\theta$: initialized TDM parameters
@@
1 for $each\ training\ step$ do
       // Collect episode with QT-TDM and add to buffer
2      $\mathcal{B}\leftarrow\mathcal{B}\cup\{s_{t},a_{t},r_{t},s_{t+1}\}_{t=0}^{T-1}$
3      for $num\ updates\ per\ episode$ do
             $\{s_{t},a_{t},r_{t},s_{t+1}\}_{t=1}^{L}\sim\mathcal{B}$;
              $\triangleright$ Sample trajectory
@@ 2      $\mathcal{B}\leftarrow\mathcal{B}\cup\{s_{t},a_{t},r_{t},s_{t+1}\}_{t=0}^{T-1}$
4            for $t=1...L$ do
                   $h_{t}=\mathit{f}_{\theta}(s\leq t,a^{1:N}\leq t)$;
                    $\triangleright$ Hidden state
                   $\hat{s}_{t+1}=\mathit{g}_{\theta}(h_{t})$;
                    $\triangleright$ Transition
                   $\hat{r}_{t}\sim\mathit{p}_{\theta}(\hat{r}_{t}|h_{t})$;
                    $\triangleright$ Reward
5
6
            $\theta\leftarrow\theta-\eta_{d}\gradient_{\theta}\mathcal{L}_{\theta}^{Dyn}$;
              $\triangleright$ Equation 6
@@
7            for $i=1...N$ do
                   $h_{t}^{i}=\mathit{f}_{\phi}(s_{t},a_{t}^{1:i-1})$;
                    $\triangleright$ Hidden state
                   $\hat{q}_{t}^{i,1:K}=\mathit{g}_{\phi}(h_{t}^{i})$;
                    $\triangleright$ Q-Values
8
            $\phi\leftarrow\phi-\eta_{q}\gradient_{\phi}\mathcal{L}_{\phi}^{Q}$;
              $\triangleright$ Equation 7
@@
9            $\bar{\phi}\leftarrow(1-\zeta)\bar{\phi}+\zeta\phi$;
10
11
Algorithm 1 QT-TDM (Training)
@@ IV-B Training
$\mathcal{L}_{\theta}^{Dyn}=\sum^{\mathit{L}}_{t=1}\Bigl{[}\mathcal{\beta}_{1}\norm{\mathit{g}_{\theta}(h_{t})-s_{t+1}}^{2}_{2}-\mathcal{\beta}_{2}\ln{\mathit{p}_{\theta}(r_{t}|h_{t})}\Bigr{]},$ (6)
@@
Require: $\theta$, $\phi$: TDM parameters, QT parameters
@@
$s_{t}$, $H$, $\mathcal{I}$: current state, len. of horizon, iterations
1
2for $n=1...\mathcal{I}$ do
3       Sample $J$ action seq. from $\mathcal{N}(\mu^{n-1},(\sigma^{n-1})^{2}\mathbf{I})$
4      Sample $J_{QT}$ action seq. using QT and TDM
      // Rollout trajectories and estimate total return $\mathcal{F}_{\varGamma}$
5      for $all\ J+J_{QT}\ action\ sequences$ do
6             for $t=0...H-1$ do
                   $h_{t}=\mathit{f}_{\theta}(s\leq t,a^{1:N}\leq t)$;
                    $\triangleright$ Hidden state
                   $\hat{s}_{t+1}=\mathit{g}_{\theta}(h_{t})$;
                    $\triangleright$ Transition
                   $\mathcal{F}_{\varGamma}=\mathcal{F}_{\varGamma}+\gamma^{t}\mathit{p}_{\theta}(\hat{r}_{t}|h_{t})$;
                    $\triangleright$ Reward
7
            // Estimate the terminal Q-value using QT and add the value of last action dim. to $\mathcal{F}_{\varGamma}$
8            $\mathcal{F}_{\varGamma}=\mathcal{F}_{\varGamma}+\gamma^{H}\max\limits_{a_{H}^{N}}Q_{\phi}(\hat{s}_{H},a_{H}^{N})$
      Update $\mu^{n}$ and $\sigma^{n}$;
        $\triangleright$ Equation 9
9
return $a_{t}\sim\mathcal{N}(\mu_{t}^{\mathcal{I}},(\sigma_{t}^{\mathcal{I}})^{2}\mathbf{I})$;
  $\triangleright$ First action is executed
10
Algorithm 2 QT-TDM (Planning)